OpenStudy (anonymous):

Write the equation for a circle that satisfies the given conditions. Center (-4,-2) and radius 14 units

OpenStudy (yrelhan4):

the equation of circle with centre x,y and radius r is (x-x1)^2 + (y-y1)^2= r^2.. good enough for a hint?

OpenStudy (anonymous):

thanks

OpenStudy (yrelhan4):

So your answer would be (x+4)^2 + (y+2)^2 = 14^2.. Click Best Response if you like my answer. :)